# Oriented Grassmannian Bundle, Normal Curvature Reduction, and Persistent Homology

### v1

@@ 1 Introduction
Persistent homology computes barcodes of filtered chain complexes constructed from a metric space $X$ [zomorodian2004computing, oudot2015persistence]. For $r>0$, the Čech complex $\check{C}(X;r)$ is the nerve of the open cover of $r$-balls $\{B_{X}(x,r)\}_{x\in X}$, and the Vietoris–Rips complex $\mathrm{VR}(X;r)$ is the abstract simplicial complex whose simplices are the finite subsets $\sigma\subset X$ with $\operatorname{diam}(\sigma)<r$. If $M$ is a Riemannian manifold, then the Čech complex $\check{C}(M;r)$ is homotopy equivalent to $M$ for $r<r_{\operatorname{conv}}(M)$ by the nerve lemma, while the systole and the filling radius bound the range of $r$ for which the Vietoris–Rips complex $\mathrm{VR}(M;r)$ has the same first and top-dimensional homology groups as $M$ [virk20201, lim2024vietoris].
The reach (or normal injectivity radius) $\mathsf{rch}_{\mathbb{R}^{D}}(M)$ of a smooth submanifold $M\subset\mathbb{R}^{D}$ is the supremum of $\tau>0$ such that every point $x\in\mathbb{R}^{D}$ with $d_{\mathbb{R}^{D}}(x,M)<\tau$ has a unique nearest point in $M$ [federer1959curvature, aamari2019estimating]. For smooth embedded submanifolds, this agrees with the supremum of radii for which the normal exponential map is injective. The reach of $M$ gives bounds on the scales $r$ where the Vietoris–Rips or Čech complex built from sufficiently dense points of $M$ in the ambient Euclidean metric [niyogi2008finding, kim2019homotopy], and the metric thickening of $M$ [adams2019metric, adams2022metric], the 1-Wasserstein analogue of the Vietoris–Rips complex, are homotopy equivalent to $M$.
@@
In both examples the indicated bottlenecks have opposite normal directions. This suggests that, for an orientable submanifold $M\subset\mathbb{R}^{D}$, embedding $M$ into an augmented ambient space, such as the oriented Grassmannian bundle $\mathbb{R}^{D}\times\mathbf{Gr}^{+}(D,d)$, can enlarge the normal injectivity radius of $M$. In the product metric on $\mathbb{R}^{D}\times\mathbf{Gr}^{+}(D,d)$, the contribution from the Grassmannian factor enlarges the lengths of curves near points of large normal curvature, so that unit-speed geodesics have smaller normal curvature, and bottleneck pairs with opposite normal directions are pushed farther apart because their oriented tangent spaces are far from each other in $\mathbf{Gr}^{+}(D,d)$.
@@
$\mathrm{vol}_{c}(M)\leq\bigl(1+c\,\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}^{2}\bigr)^{d/2}\mathrm{vol}(M),\qquad\kappa(v)>\frac{\|\nabla_{v}^{M}\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}}{\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}},$
@@
$\kappa_{c}(v)\,\mathrm{vol}_{c}(M)^{1/d}\;\leq\;\frac{\sqrt{\kappa(v)^{2}+c\|\nabla_{v}^{M}\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}^{2}}}{1+c\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}^{2}}\,\mathrm{vol}_{c}(M)^{1/d}\;<\;\kappa(v)\,\mathrm{vol}(M)^{1/d},$
@@
Similarly, in the Euclidean case we define
@@
Let $\{M_{i}\}_{i\in\mathbb{N}}$ be a sequence of smooth closed orientable $d$-dimensional submanifolds of $\mathbb{R}^{D}$ converging to a smooth closed orientable submanifold $M_{\infty}\subset\mathbb{R}^{D}$ in the Whitney $C^{1}$ topology (see Definition 2.1). Fix $c>0$. Then for every field $\mathbb{k}$ and every $j\in\mathbb{N}\cup\{0\}$,
@@ 2.1 Persistence theory
$\mathrm{VR}(X;\varepsilon_{0})\hookrightarrow\mathrm{VR}(X;\varepsilon_{1})\hookrightarrow\mathrm{VR}(X;\varepsilon_{2})\hookrightarrow\cdots$
@@
For a field $\mathbb{k}$, a persistence module (over $\mathbb{R}_{\geq 0}$) $\mathbb{V}$ is a family of vector spaces $\{V_{i}\}_{i\in\mathbb{R}_{\geq 0}}$ equipped with linear maps $v_{i}^{j}:V_{i}\to V_{j}$ for every $0\leq i\leq j$, satisfying $v_{j}^{k}\circ v_{i}^{j}=v_{i}^{k}$ whenever $i\leq j\leq k$ and $v_{i}^{i}$ is the identity map on $V_{i}$. A morphism of degree $\varepsilon$ between two persistence modules $\mathbb{V}$ and $\mathbb{W}$ is a family of linear maps $\Phi=\{\phi_{i}:V_{i}\to W_{i+\varepsilon}\}_{i\in\mathbb{R}_{\geq 0}}$ such that for all $i\leq j$,

### v2

@@ 1 Introduction
Persistent homology computes barcodes of filtered chain complexes constructed from a metric space $X$ [zomorodian2004computing, oudot2015persistence]. For $r>0$, the Čech complex $\check{C}(X;r)$ is the nerve of the open cover of $r$-balls $\{B_{X}(x,r)\}_{x\in X}$, and the Vietoris–Rips complex $\mathrm{VR}(X;r)$ is the abstract simplicial complex whose simplices are the finite subsets $\sigma\subset X$ with $\operatorname{diam}(\sigma)<r$. If $M$ is a Riemannian manifold, then the Čech complex $\check{C}(M;r)$ is homotopy equivalent to $M$ for $r<r_{\operatorname{conv}}(M)$ by the nerve lemma, while the systole and the filling radius bound the range of $r$ for which the Vietoris–Rips complex $\mathrm{VR}(M;r)$ has the same first and top-dimensional homology groups as $M$ [virk20201, lim2024vietoris, balitskiy2025geometric].
The reach (or normal injectivity radius) $\mathsf{rch}_{\mathbb{R}^{D}}(M)$ of a smooth submanifold $M\subset\mathbb{R}^{D}$ is the supremum of $\tau>0$ such that every point $x\in\mathbb{R}^{D}$ with $d_{\mathbb{R}^{D}}(x,M)<\tau$ has a unique nearest point in $M$ [federer1959curvature, aamari2019estimating]. For smooth embedded submanifolds, this agrees with the supremum of radii for which the normal exponential map is injective. The reach of $M$ gives bounds on the scales $r$ where the Vietoris–Rips or Čech complex built from sufficiently dense points of $M$ in the ambient Euclidean metric [niyogi2008finding, kim2019homotopy], and the metric thickening of $M$ [adams2019metric, adams2022metric, adams2024persistent], the 1-Wasserstein analogue of the Vietoris–Rips complex, are homotopy equivalent to $M$.
@@
In both examples, the indicated bottlenecks have opposite normal directions. This suggests that, for an orientable submanifold $M\subset\mathbb{R}^{D}$, embedding $M$ into an augmented ambient space, such as the oriented Grassmannian bundle $\mathbb{R}^{D}\times\mathbf{Gr}^{+}(D,d)$, can enlarge the normal injectivity radius of $M$. In the product metric on $\mathbb{R}^{D}\times\mathbf{Gr}^{+}(D,d)$, the contribution from the Grassmannian factor enlarges the lengths of curves near points of large normal curvature, so that unit-speed geodesics have smaller normal curvature, and bottleneck pairs with opposite normal directions are pushed farther apart because their oriented tangent spaces are far from each other in $\mathbf{Gr}^{+}(D,d)$.
@@
$\mathrm{vol}_{c}(M)\leq\bigl(1+c\,\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}^{2}\bigr)^{d/2}\mathrm{vol}(M),\qquad\kappa(v)>\frac{\|(\nabla_{v}^{M}\mathbf{II})(v,\cdot)\|_{\operatorname{HS}}}{\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}},$
@@
$\kappa_{c}(v)\,\mathrm{vol}_{c}(M)^{1/d}\;\leq\;\frac{\sqrt{\kappa(v)^{2}+c\|(\nabla_{v}^{M}\mathbf{II})(v,\cdot)\|_{\operatorname{HS}}^{2}}}{1+c\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}^{2}}\,\mathrm{vol}_{c}(M)^{1/d}\;<\;\kappa(v)\,\mathrm{vol}(M)^{1/d},$
@@
Similarly, in the Euclidean case, we define
@@
Let $\{M_{i}\}_{i\in\mathbb{N}}$ be a sequence of smooth closed orientable $d$-dimensional submanifolds of $\mathbb{R}^{D}$ converging to a smooth closed orientable submanifold $M_{\infty}\subset\mathbb{R}^{D}$ in the Whitney $C^{1}$ topology (see Definition LABEL:Whitney). Fix $c>0$. Then for every field $\mathbb{k}$ and every $j\in\mathbb{N}\cup\{0\}$,
@@ 2.1 Persistence theory
$\mathrm{VR}(X;\varepsilon_{0})\hookrightarrow\mathrm{VR}(X;\varepsilon_{1})\hookrightarrow\mathrm{VR}(X;\varepsilon_{2})\hookrightarrow\cdots,$
@@
For a field $\mathbb{k}$, a persistence module (over $\mathbb{R}_{\geq 0}$) $\mathbb{V}$ is a family of vector spaces $\{V_{i}\}_{i\in\mathbb{R}_{\geq 0}}$ equipped with linear maps $v_{i}^{j}:V_{i}\to V_{j}$ for every $0\leq i\leq j$, satisfying $v_{j}^{k}\circ v_{i}^{j}=v_{i}^{k}$ whenever $i\leq j\leq k$ and $v_{i}^{i}$ is the identity map on $V_{i}$. A morphism of degree $\varepsilon$ between two persistence modules $\mathbb{V}$ and $\mathbb{W}$ is a family of linear maps $\Phi=\{\phi_{i}:V_{i}\to W_{i+\varepsilon}\}_{i\in\mathbb{R}_{\geq 0}}$ such that for all $i\leq j$, it holds that